### v1

# Unveiling Ferroelectric-Like Behavior in Leaky Dielectrics: A Microscopic Model for Polarization Dynamics and Hysteresis Inversion

Abstract
Based on a microscopic model of nonequilibrium carrier generation in a leaky dielectric, we analytically derive hysteresis loops for the dielectric response of non-polar, non-ferroelectric materials. We demonstrate how complex dielectric responses can arise solely from energy and voltage polarity-dependent transport and asymmetries in the transfer rates. By combining Electrochemical Impedance Spectroscopy and voltammetry, we address critical questions related to the microscopic mechanisms in poorly conductive systems dominated by displacement currents. The impedance analysis, extended to higher-order harmonics, provides deeper insights into the dynamic behavior of dielectric materials, emphasizing the need to correlate impedance spectroscopy with dielectric spectroscopy for a thorough understanding of dipole relaxation and transport phenomena. Our findings offer valuable perspectives for applications in capacitors, transistors, and memory devices.
I Introduction
In the study of dielectric and polar materials, several fundamental questions are still relevant, particularly concerning the interplay between dielectric responses and carrier drift under nonequilibrium dynamics Chen et al. (2022); Wang et al. (2024), specifically in tuning dipole relaxation and leakage effects Liu et al. (2020); Wang et al. (2023). Electrochemical impedance spectroscopy (EIS) Irvine et al. (1990); Barsoukov and Macdonald (2018), dielectric spectroscopy Woodward (2021); Clark et al. (2024), and voltammetry Balke et al. (2015); Luo et al. (2023) are powerful techniques widely used in material science and nanoscale device characterization to explore these issues. However, traditional approaches to such techniques may hinder the extraction of relevant information about these nonequilibrium processes.
Polar materials such as perovskites, ferroelectrics, piezoelectrics and dielectrics in general are of significant interest due to their potential applications in sensors Cui et al. (2022), actuators Liu et al. (2023), memory devices Kittl et al. (2009), and photovoltaic cells Aqoma et al. (2024). Understanding the microscopic mechanisms in these materials, especially in poorly conductive systems dominated by displacement currents, is crucial for optimizing their performance and expanding their technological applications. This paper aims to address several critical questions, offering deeper insights into these underlying phenomena.
Firstly, we explore how the intermix of dielectric response and carrier drift affects voltammetry and impedance from a microscopic perspective. Understanding these interactions is essential to accurately interpret the data obtained from these techniques. Secondly, we investigate potential mimetic effects arising from the interplay between dielectric response and carrier drift, which might be misinterpreted. Our findings reveal that there are additional nuances to the ferroelectric mimicry explored in Ref. 15, highlighting the complexity of these interactions. Distinguishing between these effects is crucial to avoid erroneous conclusions about material properties and device behavior Esquivel-Elizondo et al. (2011); Balke et al. (2015); Vasudevan et al. (2017); Ganeshkumar et al. (2017). Furthermore, we examine the differences in parameter interpretation when using a microscopic model versus traditional equivalent circuit representations Irvine et al. (1990); Li et al. (2011); Vadhva et al. (2021); Oliveira de Lira et al. (2020); Ramírez-González et al. (2023). This comparison highlights the limitations of conventional models and the potential inaccuracies that may arise from their use.
In this context, we examine the implications of using equivalent circuits that may not fully capture the underlying microscopic processes in dielectrics. While equivalent circuits offer a valuable framework for simplifying complex systems and aiding in the interpretation of impedance data Bisquert et al. (2024); Bisquert (2024), projections into stationary circuit elements without a proper microscopic basis can sometimes result in inconsistencies or unexpected behaviors. This highlights the value of incorporating microscopic insights to enhance the accuracy and reliability of equivalent circuit models, while still benefiting from their ability to streamline device analysis and design. Additionally, we discuss how to correlate impedance analysis, dielectric spectroscopy, and higher mode generation with microscopic features and structural parameters. This correlation is vital for developing a comprehensive understanding of the material and device characteristics and we are able to provide it in fully analytical form. Lastly, we identify what might be missing in traditional impedance and voltammetry characterization, with an emphasis on multimode analysis. This approach can uncover additional dimensions of information that are often overlooked in conventional analyses.
By addressing these questions, this paper aims to enhance the reliability of EIS and voltammetry techniques, providing more precise guidelines for their application in material science and device research. This enhanced understanding is particularly crucial for systems involving dielectrics and leaky dielectrics, such as capacitors, metal-oxide interfaces in transistors, and floating gate memories. In these systems, the dynamic behavior of electric polarization plays a pivotal role, influencing performance and stability.
II Results
To analyze the effects just highlighted, we developed a model consisting of a capacitor filled with a dielectric material of size $D$, as represented in Fig. 1 (a). Inside this material, charges can be trapped in small sites of mean size $L$, which are displaced from equilibrium when subjected to an external electric field. These sites can have different shapes and size distributions that could correspond to crystallites in a polycrystalline material or grains in an amorphous one. Different dynamics of charges within the capacitor are considered in this model:
  • Charges can enter the system and be trapped in any site.
  • Charges can escape from a site due to a leakage effect and become trapped by another site.
  • Charges can escape from a site due to a leakage effect and exit the system.
  • A small amount of charges can still enter and exit the system directly without undergoing any of the previous steps.
Under an electric field, as illustrated in Fig. 1 (b) fluxes of nonequilibrium charges can produce a local polarization imbalance as mobile charges bounce within the sites following the voltage polarity. Additionally, some amount of charges can leak out from the bouncing sites, as represented in Fig. 1 (c).
The nonequilibrium charge fluctuations can be produced by localized charge traps or sources (such as a point defect or array of defects) present within the localization sites of size $L$, as they are biased at a fraction $\eta V$ of the total applied voltage $V$. In the context of a uniform distribution of sites, $\eta=L/D$ represents the inversed number of sites along a line between the two contacts. Thus, the parameter $\eta$ characterizes the voltage drop efficiency per site.
Figure 1: (a) Illustration of a capapacitor with plate separation $D$ and plate area $A$ with a dielectric medium containing charge localization sites of mean size $L$. (b) Illustration of internal fluxes of nonequilibrium charges leading to polarization fluctuations. (c) Illustration of the additional presence of leaking processes as fluxes overcoming the external barriers into the sorrounding dielectric.
Out of equilibrium, the polarization transfer function can be derived from the sum of time rates of local mobile charge generation or trapping, $\partial q_{i}/\partial t$, at a position $\langle r_{i}\rangle$ over the total volume, $\Omega=D\cdot A$, as follows
$\left.\frac{\partial P}{\partial t}\right|^{(\gamma)}=\frac{1}{\Omega}\sum_{i}\frac{\partial q_{i}}{\partial t}\langle r_{i}\rangle,$ (1)
that leads to electric dipole, $\partial q_{i}\langle r_{i}\rangle$, generation through polarization confinement ($\gamma=p$) or dipole annihilation through leaking processes ($\gamma=lk$). By assuming a uniform distribution of sites within an area $A$ the sum in Eq. 1 can be segmented over the inplane dipoles, $k$, and along the distance perpendicular to the capacitor plates, $j$, as
$\left.\frac{\partial P}{\partial t}\right|^{(\gamma)}=\frac{1}{D}\sum_{j}\frac{1}{A}\sum_{k}\frac{\partial q_{j}^{(k)}}{\partial t}\langle r_{j}\rangle,$ (2)
leading to a dependence of the polarization transfer on a combination of local charge fluxes $F^{(\gamma)}\equiv 1/A\sum_{k}\partial q_{j}^{(k)}/\partial t$.
We assume that the charge within each site redistributes either to the left or right according to the local bias, which is proportional to the total applied voltage $\eta V$, following the configuration shown in Fig. 1(a). Thus, the nonequilibrium polarization transfer described in Eq. 2 will induce a local bouncing of carriers, $\partial P/\partial t|^{(p)}$, which can be characterized in terms of carrier fluxes over left and right internal barriers. Concurrently, charges can either leak or enter the site via the external barriers, leading to a polarization fluctuation rate $\partial P/\partial t|^{(lk)}$. In a first approximation, we will assume these leaking mechanisms to be independent of the bouncing process, each with potentially distinct relaxation times, $\tau^{(p)}$ and $\tau^{(lk)}$.
Given the total number of localization sites along a line between contacts, proportional to $1/\eta$, for carriers with unit charge $q$ (where the sign may vary depending on the nature of the charge: electrons, holes, ions), the expression in Eq. 2 can be simplified to
$\left.\frac{\partial P}{\partial t}\right|^{(\gamma)}=\frac{1}{D\eta}q\left(F_{L}^{(\gamma)}L_{L}+F_{R}^{(\gamma)}L_{R}\right),$ (3)
where $F_{L}^{(\gamma)}$ and $F_{R}^{(\gamma)}$, represent the carrier fluxes to the left and right-hand side, respectively, as shown in Fig. 1 (a). And charges may accumulate (or leak) at a distance $L_{R}$, to the right, or $L_{L}$, to the left, of the generation center. This configuration also accounts for the possibility of an asymmetric energy difference between the localized charge source and the left and right sides of the site. This asymmetry can be represented by a symmetry parameter $\alpha\in[0,\infty]$, such that $\eta_{L}=\alpha\eta/(1+\alpha)$ and $\eta_{R}=\eta/(1+\alpha)$. This asymmetry can also be spatially projected, with $L_{L}=\alpha L/(1+\alpha)$ and $L_{R}=L/(1+\alpha)$, where $\alpha=1$ corresponds to the perfectly symmetric case.
In the case of electrons sources with $q=-e$ and effective mass $m^{*}$ at an effective temperature $T_{eff}$ Guarin Castro et al. (2021), the corresponding fluxes due to thermal ionization are given by
$F_{L(R)}^{(\gamma)}=\lambda^{(\gamma)}\left[1-\exp\left(\eta_{L(R)}\frac{eV}{k_{B}T_{eff}}\right)\right],$ (4)
and
$F_{R(L)}^{(\gamma)}=\lambda^{(\gamma)}\left[\exp\left(-\eta_{R(L)}\frac{eV}{k_{B}T_{eff}}\right)-1\right],$ (5)
with $\lambda^{(\gamma)}=4\pi m^{*}(k_{B}T_{eff})^{2}\exp\left(-\frac{E^{(\gamma)}}{k_{B}T_{eff}}\right)/(2\pi\hbar)^{3}$. The nonequilibrium fluxes are tuned by internal activation barriers, $E^{(p)}$, attributed to the carrier source or trapping that leads to the charge bouncing within the site. Simultaneously, an external barrier, $E^{(lk)}$, is also present and must be overcome for charge to leak out or be trapped into the site.
The eventual exhaustion of the trapping or generation process, which drives the polarization fluctuation at large voltages (independent of its polarity), can be modeled by saturating the local voltage efficiency, just for the polarization transfer and not the charge leakage. This is achieved by substituting $\eta\rightarrow\eta/\sqrt{1+\beta V^{2}}$, in Eqs. 4 and 5 for $(\gamma=p)$ introducing $\beta>0$ as an adjustable parameter so that this effect does not contribute at small amplitudes when $\beta V^{2}<<1$.
The polarization transfer functions, derived by combining Eqs. 3-5, are presented in Fig. 2 (a) for various values of the symmetry parameter $\alpha$. These functions consistently exhibit a monotonic increase, where the correlation between voltage polarity and their sign remains fixed: positive polarity corresponds to a positive sign, while negative polarity results in a negative sign. The breaking of symmetry, as reflected by varying $\alpha$, leads to an asymmetry in the function’s strength with respect to polarity. Additionally, the impact of polarization transfer exhaustion, either included ($\beta\neq 0$) or excluded ($\beta=0$), is also displayed in panel 2 (a), highlighting its influence on the overall behavior of the transfer function.
Moreover, increasing the voltage drop efficiency $\eta$ causes the apparent onset of the transfer function to extend beyond $V=0$, as depicted in Fig. 2 (b). Next, we will examine how these effects influence the tuning of electric polarization dynamics and the corresponding current-voltage characteristics.
Figure 2: (a) Contrasting transfer functions as function of applied bias for $\alpha<1$, $\alpha=1$, and $\alpha>1$ with (without) the saturation factor $\beta\neq 0$ ($\beta=0$), for $\eta=0.005$ and period $T=10\pi\tau^{(p)}$. (b) Contrasting nonequilibrium charge transfer functions as function of applied bias for varying efficiency factors $\eta^{p}$, $\alpha=1.0$, and period $T=20\pi\tau^{(p)}$. (c) Calculated polarization ($P$) as a function of the applied bias for triangular voltage sweeps, of large voltage amplitudes, in the absence of charge leakage, under parameters of the transfer function (a) and (d) calculated polarization ($P$) under parameters of the transfer function (b). The stationary results ($P_{st}$) for the polarization-voltage characteristics have been represented using red dashed curves. The respective currents (I) are provided in panels (e) and (f).
Figure 3: Small voltage amplitude polarization ($P$) in presence of both, polarization fluctuation and leakage with $\tau^{(lk)}/\tau^{(p)}=10$ with the parameters listed in Table 1 of the Appendix: (a) for a voltage sweep period, $T=2\pi\tau^{(p)}$; (b) for a voltage sweep period, $T=2\pi\tau^{(lk)}$. The stationary results ($P_{st}$) for the polarization-voltage characteristics have been represented using red dashed curves. Their respective currents (I) are displayed in panels (c) and (d).
The presence of polarization transfer, described by Eq. 3, will result in fluctuations of the polarization around certain instantaneous equilibrium values determined by the instantaneous electric susceptibility, $\chi_{\infty}$. The instantaneous polarization is given by $P_{\infty}=\varepsilon_{0}\chi_{\infty}E$, $\varepsilon_{0}$ being the vacuum permittivity, assuming a uniform electric field $E=V/D$ along the device, so the total polarization can be expressed as
$P=P_{\infty}+\delta P^{(p)}-\delta P^{(lk)}.$ (6)
Within the relaxation time, $\tau^{(\gamma)}$, approximation, these fluctuations can be described as
$\frac{d\delta P^{(\gamma)}}{dt}=-\frac{\delta P^{(\gamma)}}{\tau^{(\gamma)}}+\left.\frac{\partial P}{\partial t}\right|^{(\gamma)}.$ (7)
With these components, the displacement current can be calculated as
$j_{D}=\varepsilon_{0}\frac{dE}{dt}+\frac{dP}{dt}.$ (8)
To calculate the total current through the device, we also account for a residual carrier drift with conductance $G_{0}$ complemented by the leaking charge fluctuation $\delta n$
$I$ $=$ $j_{D}\cdot A+\left(G_{0}+\Gamma\delta n\right)V$
$=$ $C\frac{dV}{dt}+\frac{d\delta P^{(p)}}{dt}-\frac{d\delta P^{(lk)}}{dt}+\left(G_{0}+\Gamma\delta n\right)V.$
Here, $C=\varepsilon_{0}(1+\chi_{\infty})A/D$, $\Gamma=e\mu/D^{2}$, and $\mu$ represents the nonequilibrium carrier mobility de Paiva et al. (2022). The potential effects of drift current modulation with nonequilibrium charge fluctuations, $\delta n$, are detailed in Refs. Lopez-Richard et al. (2023a, b, 2024). Therefore, for the purposes of this study, we will assume $\Gamma\delta n<<G_{0}$, maintaining just a residual leaking conductance, $G_{0}\neq 0$, in addition to the displacement contributions. For simplicity, we have set $G_{0}=1$ as the unit of conductance in the presence of leakage.
Figure 2 (c) illustrates an example of hysteresis in the polarization-voltage loop under a periodic triangular input, calculated according to Eq.6 for a symmetrical transfer function with $\alpha=1$ and in the absence of leakage. The arrows indicate the direction of the loop. Notably, the generated polarization resembles a ferroelectric response Kumar et al. (2021); Sekine et al. (2022); Manzi et al. (2023). The polarization is not fully saturated due to the linear voltage dependence assumed for $P_{\infty}=\varepsilon_{0}\chi_{\infty}V/D$. When increasing the parameter $\eta$, under a symmetric transfer function, the voltage onset that triggers polarization fluctuation increases, resulting in an apparent pinched polarization hysteresis, as shown in panel 2 (d). This resembles the traces of a double hysteresis loop, which are typically associated with antiferroelectric behavior Pintilie et al. (2008); Wang et al. (2020); Luo et al. (2023). In this scenario, the apparent double hysteresis loops emerge due to the widening voltage onset of the transfer function. The stationary result for the polarization-voltage characteristics, $P_{st}$, obtained by setting $d\delta P^{(p)}/dt=0$ in Eq. 6 has been added to both Figs 2 (c) and (d) as reference by using dashed curves.
The corresponding current-voltage loops, calculated according to Eq. II, are depicted in panels 2 (e) and (f), respectively. The opening of the hysteresis response depends not only on the relative magnitude of nonequilibrium contributions ($\delta P$ in this case) compared to the stationary values, but also on the period of the cyclic voltage. Optimal conditions for a memory response, when the voltage period approaches $T\sim 2\pi\tau^{(\gamma)}$, are discussed in Refs. 26; 35; 27.
For small voltage amplitudes, the saturation effects can be neglected ($\beta V^{2}\approx 0$, after Eqs. 4 and 5) and solving Eqs. 3-II in the presence of leakage, one may obtain results as those depicted in Figs 3 (a)-(d). In order to contrast the potential time-scale difference we have set $\tau^{(p)}=\tau^{(lk)}/10$ with symmetric polarization transfer ($\alpha=1$). In Fig. 3 (a), a ferroelectric-like counterclockwise loop is achieved by setting the voltage sweeping period to $T=2\pi\tau^{(p)}$, which maximizes polarization fluctuations while keeping the leakage contribution negligible. By adjusting the voltage sweep period to $T=2\pi\tau^{(lk)}$, the polarization loop direction can reverse, as shown in Fig. 3 (b). Under this condition, the voltage sweep aligns with the charge leakage dynamics, enabling the hysteresis loop to switch direction. Experimentally, clockwise reversible polarization loops have been observed in Ref. 36, and similar effects have been investigated in Ref. 37. In addition to frequency tuning, the balance between polarization transfer and charge leakage is significantly influenced by temperature, governed by the parameters $\lambda^{(p)}$ and $\lambda^{(lk)}$ as defined after Eq. 5, as confirmed in Ref. 36. The impact of temperature tuning on these effects will be further addressed in the discussion.
Note that despite the reversal of the polarization loop, the current loops have consistently maintained a clockwise direction throughout both Fig. 2 and Fig. 3. This behavior arises from the relatively strong contribution of the capacitive term in Eq. II, $CdV/dt$. For a negligible capacitive contribution (with $C\rightarrow 0$), a reversal of the current loops is also expected. In Fig. 4, we compare two scenarios with parameters minimizing the capacitive term: the left column shows a larger value of $\lambda^{(p)}$, where no loop reversal is observed within the analyzed period range, while for a smaller $\lambda^{(p)}$, a counterclockwise current loop emerges as the period increases. This effect has been experimentally observed in Ref. 37 and attributed to the detrapping of electrons from interface traps, in agreement with the theoretical mechanisms described in our model.
Figure 4: Left column: current-voltage loops using the same parameters as in Fig. 3 (a) as listed in Table 1 of the Appendix, but with capacitance $C\rightarrow 0$, varying the period, $T$. Right column: current-voltage loops with the same parameters but $\lambda^{(p)}=50.0$, varying the period, $T$.
Given the period tunability of the polarization and displacement response, a spectroscopic analysis becomes paramount for a comprehensive characterization of the transport behavior. By analyzing the frequency-dependent response, we can better understand how charge dynamics, polarization transfer, and leakage processes interact across different time scales. This allows us to probe the underlying mechanisms in greater detail and assess the material’s overall dielectric and transport properties more effectively.
Under a sinusoidal voltage input, $V=V_{0}\cos(\omega t)$, and by expanding Eqs. 3-5 in powers of $\eta eV_{0}/(k_{B}T_{eff})$ the solution of Eq. 7 can be derived analytically. The stable response, once all transient contributions have dissipated, can be expressed as a combination of harmonics Lopez-Richard et al. (2023b)
$I=V_{0}\sum_{n=0}\left[G_{\left(n\right)}\cos\left(n\omega t\right)-B_{\left(n\right)}\sin\left(n\omega t\right)\right].$ (10)
Thus, assuming $\eta eV_{0}/(k_{B}T_{eff})<<1$ the $n-mode$ conductance $G_{(n)}$ and susceptance $B_{(n)}$, up to second order, are given by
$G_{(1)}=\frac{A}{V_{0}}\left\{\frac{\sigma_{1}^{(p)}\left[\omega\tau^{(p)}\right]^{2}}{1+\left[\omega\tau^{(p)}\right]^{2}}-\frac{\sigma_{1}^{(lk)}\left[\omega\tau^{(lk)}\right]^{2}}{1+\left[\omega\tau^{(lk)}\right]^{2}}\right\}+G_{0},$ (11)
Figure 5: Apparent circuit representations for: (a) the first and (b) the second mode of the dielectric impedance. First (c) and second (d) order mode Nyquist plots for: $G_{0}=1.0$ and $C=0.01$. The rest of the parameters is listed in Table 1 of the Appendix. The arrows point towards the frequency growth direction.
$B_{(1)}=\frac{A}{V_{0}}\left\{\frac{\sigma_{1}^{(p)}\omega\tau^{(p)}}{1+\left[\omega\tau^{(p)}\right]^{2}}-\frac{\sigma_{1}^{(lk)}\omega\tau^{(lk)}}{1+\left[\omega\tau^{(lk)}\right]^{2}}\right\}+\omega C,$ (12)
and
$G_{(2)}=\frac{2A}{V_{0}}\left\{\frac{\sigma_{2}^{(p)}\left[\omega\tau^{(p)}\right]^{2}}{1+\left[2\omega\tau^{(p)}\right]^{2}}-\frac{\sigma_{2}^{(lk)}\left[\omega\tau^{(lk)}\right]^{2}}{1+\left[2\omega\tau^{(lk)}\right]^{2}}\right\},$ (13)
$B_{(2)}=\frac{A}{V_{0}}\left\{\frac{\sigma_{2}^{(p)}\omega\tau^{(p)}}{1+\left[2\omega\tau^{(p)}\right]^{2}}-\frac{\sigma_{2}^{(lk)}\omega\tau^{(lk)}}{1+\left[2\omega\tau^{(lk)}\right]^{2}}\right\},$ (14)
where
$\sigma_{1}^{(\gamma)}=\lambda^{(\gamma)}\frac{\left[\alpha^{2}+1\right]}{\left[1+\alpha\right]^{2}}\frac{\eta eV_{0}}{k_{B}T_{eff}},$ (15)
and
$\sigma_{2}^{(\gamma)}=\frac{\lambda^{(\gamma)}}{2}\frac{\left[\alpha^{3}-1\right]}{\left[1+\alpha\right]^{3}}\left[\frac{\eta eV_{0}}{k_{B}T_{eff}}\right]^{2}.$ (16)
Note that, for $n=1$ the charge bouncing contribution, proportional to $\sigma^{(p)}_{1}$, is positive defined while the leakage term, proportional to $\sigma^{(lk)}_{1}$, is negative. This will be the cause of the apparent inductive traces in the fundamental mode if $B_{(1)}<0$.
In this representation, the impedance per mode can be expressed as Lopez-Richard et al. (2023b)
$Z_{(n)}=\frac{G_{(n)}}{[G_{(n)}]^{2}+[B_{(n)}]^{2}}-i\frac{B_{(n)}}{[G_{(n)}]^{2}+[B_{(n)}]^{2}}.$ (17)
This formulation provides an impedance spectroscopic characterization that extends beyond the traditional fundamental mode, Wang et al. (2021); Lazanas and Prodromidis (2023) offering deeper insights into the system’s behavior.
Figure 6: First mode contributions to the real (a) and imaginary (b) part, of the dielectric permittivity, and the corresponding second-order contributions to the real (c) and (d), and imaginary parts for the set of parameters used in Fig. 5 (c) and (d) as listed in Table 1 of the Appendix.
According to the frequency dependence described by Eqs. 11- 14, the apparent circuits for the first and second mode impedance is depicted in Figs. 5 (a) and (b), respectively. Note that, for a resistance, $R^{\prime}$, in series with a capacitor, $C^{\prime}$ the effective admittance is
$Y=\frac{1}{R^{\prime}}\left[\frac{\left(\omega\tau^{\prime}\right)^{2}}{1+\left(\omega\tau^{\prime}\right)^{2}}+i\frac{\omega\tau^{\prime}}{1+\left(\omega\tau^{\prime}\right)^{2}}\right],$ (18)
with $\tau^{\prime}=R^{\prime}C^{\prime}$. Thus, following Eqs. 11 and 12, the apparent resistances and capacitances in panel 5 (a) are defined as follows: $R\equiv 1/G_{0}$, $R^{(p)}_{1}=V_{0}/(A\sigma^{(p)}_{1})$, and $C^{(p)}_{1}=\tau^{(p)}/R^{(p)}_{1}$, while $R^{(lk)}_{1}=-V_{0}/(A\sigma^{(lk)}_{1})$, and $C^{(lk)}_{1}=-\tau^{(lk)}/R^{(lk)}_{1}$. It is important to note that, within this representation, both the resistive and reactive components of the leakage branch will have negative signs, regardless of the combination of parameters in the transfer function. In turn, for the second order term, the apparent circuit is represented in panel 5 (b). The resistances and capacitances in this case, according to Eqs. 13 and 14, are defined as $R^{(p)}_{2}=2V_{0}/(A\sigma^{(p)}_{2})$, and $C^{(p)}_{2}=2\tau^{(p)}/R^{(p)}_{2}$, while $R^{(lk)}_{1}=-2V_{0}/(A\sigma^{(lk)}_{2})$, and $C^{(lk)}_{2}=-2\tau^{(lk)}/R^{(lk)}_{2}$. In this case, the ambiguity in the values of $\alpha\neq 1$ in the presence of asymmetry prevents a strict sign definition for $\sigma^{(\gamma)}_{2}$ in Eq. 16 and, consequently, for the sign of the apparent resistance or capacitance of the second mode branch.
Two examples of multimode impedance maps for the first and second order modes have been displayed in Figs. 5 (c) and (d), respectively. Finite values of the residual conductance $G_{0}$ and the geometric capacitance, $C$, have been considered assuming $\tau^{(lk)}=10\tau^{(p)}$. For $n=1$, at low frequencies, an apparent inductive loop emerges, as shown by the positive values of $\text{Im}(Z_{(1)})$ in panel 5(c). These effects are often referred to as negative capacitances Cheng et al. (2019), though in this case, they arise from the interplay between opposing contributions in the system. Specifically, the opposing signs of the charge storage (bouncing) and leakage terms in Eq. 12 lead to this behavior. This dynamic gives rise to inductive-like responses without invoking actual inductance, manifesting as apparent negative capacitance in the impedance spectrum.
Figure 7: First mode dielectric permittivity as a function of temperature, $T_{eff}$ for increasing AC voltage period and the parameters listed in Table 1 of the Appendix.
The contribution of the second mode, shown in panel  5(d), exhibits a noticeable transition in the impedance character as the frequency increases, shifting from the second to the fourth quadrant. This behavior arises as both the real and imaginary components of the impedance reverse sign. The transition is driven by an evolving imbalance between the opposing contributions of charge leakage and charge bouncing. As frequency increases, the dominance of one mechanism over the other shifts, leading to this distinctive impedance response. This highlights the complex interplay between dynamic processes that cannot be captured by simpler, first-order approximations.
In analyzing the transport response of a dielectric, it is important to note the correlation between impedance spectroscopy and dielectric spectroscopy. While impedance spectroscopy provides insights into the resistive (real part) and reactive (imaginary part) components of impedance, dielectric spectroscopy measures the dielectric permittivity and loss as a function of frequency, offering detailed information about dipole relaxation processes and polarization mechanisms within the material. In our model, these techniques are connected through the multimode expression for the electric susceptibility in nonlinear systems
$P=P_{\infty}+\varepsilon_{0}\sum_{n=1}\chi_{(n)}E^{n}.$ (19)
Thus, by linking this expansion with Eqs. II and 10, we can generalized the permittivity definition in this multimode perspective (including the residual conductance) as
$\varepsilon_{(n)}=\varepsilon_{0}\chi_{(n)}+\delta_{n,1}\left[\varepsilon_{0}(1+\chi_{\infty})-i\frac{G_{0}}{\omega}\frac{D}{A}\right].$ (20)
The current in Eq, II can be thus redefined in the complex plane as
$I=A\sum_{n=1}i\varepsilon_{(n)}n\omega\left(\frac{V_{0}}{D}\right)^{n}e^{in\omega t},$ (21)
that, according to Eq. 10 leads to the following correlation between the permittivity, the conductance, and susceptance per mode
$\varepsilon_{(n)}=\frac{[B_{(n)}-iG_{(n)}]}{n\omega}\frac{D^{n}}{AV_{0}^{n-1}},$ (22)
for $n=1,2,...$ The real and imaginary contributions to the first and second-order terms of the electric permittivity, corresponding to the parameters in Figs. 5 (c) and (d), are depicted in Fig.6. The behavior expected for the real part of the $n=1$ permittivity is $\text{Re}[\varepsilon_{(1)}]\propto B_{(1)}/\omega$ in Eq. 22. This results in step-up or step-down functions as a function of frequency (in logarithmic scale) depending on the leak or polarization contributions, respectively, as shown in Fig. 6 (a). Due to the presence of residual conductance, the corresponding imaginary part in panel 6 (b) diverges for decreasing $\omega$ since $\text{Im}[\varepsilon_{(1)}]\propto-G_{0}/\omega$ in Eq. 20. For the second mode, step functions (in logarithmic scale) are also expected for $\text{Re}[\varepsilon_{(2)}]\propto B_{(2)}/\omega$ for the same reasons as for the first mode, as depicted in Fig. 6 (c). The imaginary parts in this case collapse at low and high frequencies and peak at $\omega=1/(2\tau^{(lk)})$.
As discussed earlier, the relative contributions of polarization transfer and charge leakage (captured by the parameters $\lambda^{(\gamma)}$ after Eq. 5) can be modulated by varying the temperature. This temperature-dependent tuning of the dielectric response, combined with frequency adjustments, offers valuable insights into the dynamic behavior of the dielectric material. To analyze this further, we can utilize the explicit expression for the first-mode permittivity according to Eqs. 22 and 12
$\text{Re}\left[\varepsilon_{(1)}\right]=\varepsilon_{0}(1+\chi_{\infty})+\frac{4\pi m^{*}De}{(2\pi\hbar)^{3}}k_{B}T_{eff}\frac{(1+\alpha^{2})}{(1+\alpha)^{2}}\eta\left[\frac{\tau^{(p)}}{1+\left[\omega\tau^{(p)}\right]^{2}}e^{-\frac{E^{(p)}}{k_{B}T_{eff}}}-\frac{\tau^{(lk)}}{1+\left[\omega\tau^{(lk)}\right]^{2}}e^{-\frac{E^{(lk)}}{k_{B}T_{eff}}}\right]$ (23)
that reveals a peculiar temperature tuning behavior in presence of leakage. By establishing contrasting relaxation times for polarization fluctuation and charge leakage, $\tau^{(p)}<\tau^{(lk)}$, along with different values for the potential barriers, $E^{(p)}<E^{(lk)}$, the temperature dependence of this function can shift. As displayed in Fig. 7, at high frequencies $\omega$ (low period $T=2\pi/\omega$) it exhibits a monotonic trend, but as the frequency decreases, it transforms into a non-monotonic pattern resembling a relaxor response Ramesh et al. (2016); Zhou et al. (2024). At low temperatures, the dielectric tuning of polarization transfer dominates, while at higher temperatures, the leakage contribution, which reduces the relative permittivity as described in Eq. 23, becomes more significant.
III Conclusions
In this work, we have developed a comprehensive model to describe the dielectric response of non-polar, non-ferroelectric materials, with a particular focus on the interplay between polarization transfer and charge leakage mechanisms. By employing a microscopic approach, we derived analytical expressions for the polarization and current dynamics under cyclic voltammetry and impedance spectroscopy, which allow us to probe the complex transport behavior in leaky dielectrics. Our model successfully reproduces key experimental phenomena, including the inversion of polarization loops, as observed in experimental studies, such as those reported in ferroelectric-like materials. This polarization loop inversion, driven by changes in the voltage sweep period and the dominant contribution of either polarization transfer or leakage, underscores the tunability of the dielectric response.
Furthermore, the model demonstrates how ferroelectric-like hysteresis loops can emerge in non-ferroelectric materials purely due to the presence of non-equilibrium dipoles and energy-dependent transfer rates, providing new insights into the mimicry of ferroelectric traces induced by the presence of defects. Our findings closely match experimental observations of clockwise and counterclockwise reversible polarization loops, which are often misinterpreted as evidence of ferroelectricity. By clearly distinguishing these effects from true ferroelectric behavior, our study not only offers a theoretical foundation for such experimental results but also contributes to the accurate characterization of materials used in capacitors, transistors, and memory devices.
The model’s flexibility in incorporating higher-order harmonics and its extension to multimode impedance analysis further enrich the understanding of dielectric systems dominated by displacement currents. Ultimately, this work provides a valuable framework for interpreting complex dielectric responses, paving the way for more precise experimental characterization and improved device design in systems where distinguishing between real and apparent ferroelectric behavior is crucial.
IV Appendix
Figure $\tau^{(lk)}/\tau^{(p)}$ $\lambda^{(lk)}/\lambda^{(p)}$ $\eta$ $\alpha$
2(a) —–/1.0 —– $0.005$ —–
2(b) —–/1.0 —– $\sim$ 0.5 1.0
3(a) 1.0/0.1 10.0/100.0 $0.01$ 1.0
3(b) 1.0/0.1 10.0/100.0 $0.01$ 1.0
4(a) 1.0/0.1 10.0/100.0 $0.01$ 1.0
4(e) 1.0/0.1 10.0/50.0 $0.01$ 1.0
5 1.0/0.1 1.0/2.0 $0.01$ 10.0
6 1.0/0.1 1.0/2.0 $0.01$ 10.0
7 1.0/0.1 —– $0.01$ 1.0
Table 1: Parameters used in each figure.
Acknowledgements.
This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil (CAPES) and the Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brazil (CNPq) Proj. 311536/2022-0.